# DID Connect: Authentication in TLS with Decentralized Identifiers and Verifiable Credentials

### v1

@@ Abstract
Authentication in TLS is predominately carried out with X.509 digital certificates issued by certificate authorities (CA). The centralized nature of current public key infrastructures, however, comes along with severe risks, such as single points of failure and susceptibility to cyber-attacks, potentially undermining the security and trustworthiness of the entire system. With Decentralized Identifiers (DID) alongside distributed ledger technology, it becomes technically feasible to prove ownership of a unique identifier without requiring an attestation of the proof’s public key by a centralized and therefore vulnerable CA. This article presents DID Connect, a novel authentication scheme for TLS 1.3 that empowers entities to authenticate in a TLS-compliant way with self-issued X.509 certificates that are equipped with ledger-anchored DIDs instead of CA-issued identifiers. It facilitates the exchange of tamper-proof and 3rd-party attested claims in the form of DID-bound Verifiable Credentials after the TLS handshake to complete the authentication with a full identification of the communication partner. A prototypical implementation shows comparable TLS handshake durations of DID Connect if verification material is cached and reasonable prolongations if it is obtained from a ledger. The significant speed improvement of the resulting TLS channel over a widely used, DID-based alternative transport protocol on the application layer demonstrates the potential of DID Connect to become a viable solution for the establishment of secure and trustful end-to-end communication links with decentrally managed digital identities.
@@ I Introduction
This article presents DID Connect, a novel authentication scheme for TLS that empowers both parties to alternatively authenticate either with pseudo-anonymous DIDs or in combination with VCs instead of using CA-issued X.509 certificates or raw public keys. With DID Connect, self-issued X.509 certificates are used by the involved parties as a container to communicate their DIDs and associated verification material without violating the X.509v3 specification. This makes it possible with DID Connect to rely on the established way parties are authenticated during the TLS handshake by only - in addition - integrating a custom certificate validation process tailored to DID-equipped X.509 certificates. Atop the pseudo-anonymous authentication during the TLS handshake with DIDs, in DID Connect, the parties can additionally exchange DID-bound VCs afterwards, to complete the authentication scheme with a clear identification of subjects that are referred to by the pseudo-anonymous DIDs used during the TLS handshake. Since DID Connect remains fully compliant with TLS 1.3, the parties establishing a TLS connection with DID Connect benefit from the widespread support, level of standardization, and thoroughly investigated security of TLS with X.509-based authentication and, at the same time, take advantage of the key properties of decentralized identity management, namely increased data privacy and protection, improved robustness against failures, and enhanced security due to a lack of a single point of vulnerability.
@@ V Concept
With DID Connect, each entity - the client and the server - is assumed to be equipped with its own DID. The corresponding DID docs are anchored in a commonly operated and governed distributed ledger, accessible by all entities. The pseudo-anonymous authentication of an entity occurs within the TLS 1.3 handshake. Subsequently, the identification of the subject takes place at the new identification sub layer via the exchange of VCs immediately following the successful completion of the TLS handshake. The identification step is optional because both parties might know each other and each others’ DIDs from past interactions. Similar to the approach presented in [17], self-signed X.509 certificates are used in DID Connect during the TLS handshake as a standardized means to communicate the identifier as well as the public key of the certificate’s subject. In contrast to [17], a W3C DID is used as an identifier. It is stored in the SAN extension instead of the subject field, as sketched in [19], due to the strict naming rules of the X.509 subject field and the 64 byte limit of the common name (CN) attribute. The public key of the self-signed X.509 certificate corresponds to a public key of the subject’s DID doc in the distributed ledger. An exemplary message exchange of DID Connect, comprising the pseudo-anonymous authentication with DIDs in the TLS handshake and the identification with VCs at the application layer, is illustrated in Fig. 4.
@@
Figure 4: Using DIDs in X.509 certificates during the TLS handshake for the pseudo-anonymous authentication and VPs afterwards to (mutually) identify entities after the TCP connection establishment with DID Connect.
@@
As described in Section II, the client and server verify each others certificate, the CertificateVerify and the Finished messages. With a CA-issued X.509 certificate, the integrity needs to be tested by verifying the digital signature of the CA. Moreover, if intermediary CAs are present then the chain of certificates needs to be verified all way up to the root CA. However, if a self-signed X.509 certificate’s subject refers to a DID, the certificate validation becomes significantly simpler. The integrity of a self-signed and DID-based X.509 certificate is a prerequisite, but does not get explicitly tested through the verification of the certificate’s digital signature. DID Connect uses the certificate only as a vehicle to communicate the binding. However, the validity of the binding between the given DID in the SAN field and the public key in the Subject Public Key Info field needs to be checked with other means. Since the distributed ledger guarantees the uniqueness of a DID and that only the DID owner is able to alter the corresponding DID doc, it can, in the role of an independent technical trust anchor, confirm or deny the validity of a binding between a DID and a public key. If a binding’s validity is confirmed by or through the distributed ledger, the verifying entity still does not know whether the other entity is the DID owner or an entity just pretending to own the DID. The DID ownership proof is missing. Since the CertificateVerify message is signed with the private key corresponding to the public key of the certificate, it serves as a valid DID ownership proof. Technically. the self-signed X.509 certificate would actually only need to contain the DID, since the public key could be queried by resolving the DID doc from the ledger. However, providing the public key up front within the certificate speeds up the validation process because the verification of the CertificateVerify message can be carried out in parallel to the binding check via the ledger. The order of validation checks is therefore irrelevant. If the binding and the CertifyVerify message are both valid, the pseudo-anonymous authentication with a DID is completed.
@@
With DID Connect, the entities are free to choose a suitable form of authentication based on the scenario-specific requirements, the presence of CA-issued X.509 certificates and/or DIDs/VCs, and potentially prior knowledge about each other. DID Connect supports both, one-way and mutual authentication. It gives entities the opportunity to fall back on CA-issued X.509 certificates for the authentication in case self-signed X.509 certificates with DIDs are not supported. An entity can either authenticate with a CA-issued X.509 certificate, a self-issued and DID-based X.509 certificate only, or in combination with a VC. In hybrid mode, one party can make use of a self-issued and DID-based X.509 certificate or in combination with a VC to authenticate itself while the other party falls back on a CA-issued X.509 certificate. If the client does not demand the server to use a specific identifier for the authentication (by using the SNI extension), the server can provide a default CA-based X.509 certificate (Cerdef) or a default DID (DIDdef) within a self-issued X.509 certificate. Table I lists the numerous authentication options the entities can choose from together with the extensions present in the ClientHello and ServerHello messages. Since the extensions CPP and SPA consistently co-occur whenever at least one party expresses an intention to showcase VCs, they have been excluded from the table for the sake of readability.
@@ VI Evaluation
DID Connect represents an alternative authentication scheme for TLS 1.3, bridging the gap from authentication with centrally-approved to ledger-approved identification material. The practicability of DID Connect depends upon how it performs in terms of speed against the conventional way of authentication in TLS with CA-issued X.509 certificates. However, in contrast to the legacy authentication scheme, DID Connect comes in two flavors: a pseudo-anonymous authentication with DIDs only or an identification with VCs in addition. A meaningful assessment of its performance is therefore only possible by comparing both variants independently with their contemporary counterparts. The practical usefulness of DID Connect also hinges on the bandwidth utilization of the resulting TLS channel. Hence, a comparison against an alternative data transport solution that also employs DIDs for authentication purposes is required to properly evaluate the proposed solution in terms of efficiency.
@@ VI-A Methodology
From a conceptual point of view, a self-issued X.509 certificate with a subject of choice in the subject field is the closest analog to a self-issued x.509 certificate with a DID in the ASN field. In both cases, the certificate owner can prove ownership of the contained public key. Both types of certificates lack a proof of the binding’s validity by a 3rd party. However, with VDR-anchored DIDs, this proof is provided by the VDR. With legacy self-issued X.509 certificates, there is no such standardized and trustful out-of-bound option to check the binding’s validity. Hence, its validity can’t be tested unless the identifier results from the key itself, e.g., through hashing as applied in the Bitcoin network to derive the wallet address from the public key. A derived identifier is of pseudo-anonymous nature like a DID. It does not share the property of being resolvable but its uniqueness is guaranteed probabilistically in the same way as with DIDs. The pseudo-anonymous authentication scheme with DIDs in DID Connect was therefore evaluated by testing its performance against the authentication by means of a self-issued X.509 certificate with a derived identifier. If the pseudo-anonymous authentication is supplemented by an identification with VCs afterwards then DID Connect accomplishes the same goal as an authentication with legacy X.509 certificates issued by CAs, namely proving to own a 3rd-party issued identifier. First, the DID ownership is proven during the TLS handshake with self-issued and DID-equipped X.509 certificates. Afterwards, the subject referred to by the DID is revealed with a DID-bound VC that contains the 3rd party attested identifier and optionally additional identity claims of the subject. The combination of a pseudo-anonymous authentication and an identification with DID Connect is therefore evaluated by comparing it as a whole against the conventional authentication with CA-issued X.509 certificates.
@@
For the evaluation of the bandwidth utilization, a TLS channel established with DID Connect was compared against a communication link that uses DIDComm [27] instead. DIDComm is an application layer transport protocol built atop the concept of DIDs. It enables two parties to exchange messages in form of JSON Web Encryption (JWE) envelopes that are encrypted using the verification material found within each others’ DID docs, ensuring end-to-end confidentiality, integrity, and authenticity. DIDComm operates without the need for established connections and is agnostic to the transport layer protocol, with implementations available across various programming languages. Despite being connectionless and not being a transport layer protocol, it was selected for the comparison due to its widespread use, the ongoing standardization by the DIF, and the lack of an available DID-enabled transport layer protocol.
@@ VI-B Setup
The proof of concept, comprising a client and a server, was implemented with Python 3.10 using OpenSSL 3.0.2 wrapped with pyOpenSSL 23.3. The VC presentation protocol was built from scratch, in compliance with the DIF Presentation Exchange 2.X.X specification. The VCs were encoded as SD-JWTs because of its simplicity compared to JSON-LD and other formats. For the performance comparison, the prototype relies on Veramo’s implementation of DIDComm v2444https://veramo.io/. The distributed ledgers used for the experiments were powered by Hyperledger Indy [^5] https://www.hyperledger.org/projects/hyperledger-indy. During the experiments, DID Connect was tested with different ledger realizations. A locally deployed ledger based on the VON Network project [^6] https://github.com/bcgov/von-network was used to simulate a setup in which the client and server are each assumed to be co-located with a node of the ledger, making a ledger access practically a local rather than a remote operation. To simulate remote ledger access, two publicly accessible ledgers were used: the test ledger of the BCovrin Test Indy Network operated by the Province of British Columbia [^7] http://test.bcovrin.vonx.io/ and the test ledger of the German IDunion project [^8] https://idunion.org/. The latter is commonly governed and operated by the IDunion project members, coming from various economic sectors such finance, transportation, telecommunication, and manufacturing. In the experimental setup, the client was operated on a machine with an Intel Core i5-6200U CPU with 2.30 GHz and 8 GB RAM while the server resided on a machine with an Intel Core i7-4790K CPU with 4.00 GHz and 32 GB RAM. Ubuntu-based KDE Neon 5.27 served as the operating system on both machines. They were located 7 km apart with an average ping time of 12.61ms and a bandwidth of 52 Mb/s (download) and 11 Mb/s (upload) of the client and respectively 57 Mb/s (download) and 22 Mb/s (upload) of the server. Access to the Internet was provisioned by Wi-Fi access points on both sites.
@@ VI-C Results
The results show that if VDR-anchored DIDs are cached at the client and server (config. V) or peer DIDs are used instead (config. VI), the TLS handshake with DID Connect performs similar in terms of speed as their legacy counterpart (config. I). As soon as a DID doc resolution via a VDR is required, the TLS handshake with DID Connect takes significantly longer. For example, a resolution via a locally deployed VON ledger at the server in config. VIII increases the total handshake time in comparison to config. VI by a factor of 3.23 while it increases by factor of 31.58 with the BCoverin ledger in config. X and 12.12 with the IDunion ledger in config. XII. The reasons for the high resolution times via a remote ledger could be manifold, ranging from the implementation of the resolution logic and the ledger itself, over the communication latency with the ledger and the amount of transactions in the ledger, to the genesis file that specifies the settings for the ledger access. The prototype was operated in Germany, with a low latency towards the Germany-based IDunion ledger and a higher latency towards the BCoverin ledger operated in western Canada. The BCoverin ledger persisted approx. 640k transactions while the IDunion test ledger contained only approx. 117k transactions. During the evaluation, the Indy CLI driver [^9] https://github.com/hyperledger/indy-cli-rs provided by the Hyperledger Indy project was used to resolve the DID docs as it turned out in prior experiments with a locally deployed VON ledger that the resolution via a locally deployed Universal Resolver [^10] https://dev.uniresolver.io/ or directly with the Indy driver of the Universal Resolver is on average 2.03 respectively 1.89 times slower than a resolution with the Indy CLI. A genesis file points to a set of ledger nodes usable for the resolution. This selection is of utmost importance since slow or non-reachable ledger nodes in the genesis file negatively impact the overall resolution performance, as experienced with different genesis file versions for the IDunion ledger.
@@
In a second experiment, the durations of the mutual identification step after the TLS handshake were measured in addition for the configs. VI to XII. Outliers (approx. 1%) were removed since non-reachable ledger nodes during the DID resolution led to timeouts that significantly skewed the average durations. Since the pseudo-anonymous authentication together with the identification step of DID Connect reassembles an authentication with a CA-issued X.509 certificate, the durations for the latter were measured in configs. II to IV for comparison reasons. The messages exchanged by the presentation protocol during the identification in the proof of concept differs slightly to the sequence of messages sketched schematically in Fig. 4. The DIF Presentation Exchange specification demands that a message containing the VP is the response to a verifier’s request to present a VP. Hence, in the prototype, each party triggers the other party’s presentation of a VP by requesting it explicitly. However, the mutual VP presentations in the prototype were conducted in parallel to improve the identification step’s performance.
@@
The results show that when compared with an authentication based on CA-issued X.509 certificates, it takes at the client about 3.19 and 6.43 times longer to identify a subject with DID Connect in configs. VI and VII and about 42.94 and 34.86 times longer if a remote ledger is queried for the DID docs in configs. X and XII. The times to resolve the DID docs of the client, server, and the issuers of the VCs dominate the overall time of the pseudo-anonymous authentication and identification steps of DID Connect, making the DID resolution the major bottleneck of the proposed solution. In all configs. from VI to XII, it seems as it takes significantly longer for the client to identify the server than vice versa. This phenomenon results from the fact that the client has to wait for the server to process the client’s Certificate, CertificateVerify, and Finished messages before the server can proceed with the identification step. Since the client emits the first identification message, namely the VC presentation request, immediately after emitting the Finished message, it is already engaged in the identification while the server is still busy in the TLS handshake. However, an identification step with VCs is assumed to happen only once two entities have to become known to each other. Once this is done, a pseudo-anonymous authentication with DIDs is sufficient since the subjects the DIDs refer to are known. If commonly agreed upon DID doc update policies are established upon all participants of a ledger, DID docs might only change at regular or defined points in time so that most of the time locally cached versions of DID docs are sufficient for a secure TLS connection establishment. A pseudo-anonymous authentication in config. VI with DID Connect is by a factor of 1.05 even slightly faster compared to the authentication in config. I. In contrast to DID Connect and due to tight coupling of a CA-issued identifier and the corresponding key material in a signed artifact, the authentication with CA-issued X.509 certificates needs to be repeated at every TLS connection establishment.
@@
TABLE II: Kilobytes transferred with DID Connect ($\Diamond$) or DIDComm ($\blacklozenge$) depending on number of packets per session and payload size
Finally, the performances of TLS 1.3 (with DID Connect) was compared with DIDComm version 2 in terms of bandwidth utilization and the average time to transfer data in the experimental setup. For a fair comparison, DIDComm applied TCP as the transport layer protocol. The total bytes transferred by the client towards the server were measured for different payload sizes and number of packets per session. It includes two session tickets for TLS with DID Connect of total size 1.5 kB for a potential session resumption. The results are shown in Table II. For the data transfer time measurements, the client established a TLS connection (with DID Connect) to the server and transferred a certain amount of packets via the same session. A TCP session was established instead of TLS for the measurements with the DIDComm protocol. The average time for the client to send a bulk of packets to the server in the experimental setup depending on the amount of packets per bulk after 50 runs are shown in Fig. 6.
For payload sizes up until 1 kB, the initial protocol overhead with DID Connect for the transfer of a single packet is significantly higher than the one experienced with DIDComm because of the required TLS handshake. However, once the session is established, the protocol overhead with DID Connect/ TLS becomes smaller in comparison to DIDComm. The point at which a connection with DIDConnect utilizes the channel better depends on the payload size. It needs 5 packets with payload size of 1 to 100 bytes for a TLS connection to be more efficient than DIDComm or 4 packets for payload size 1 kB. Since DIDComm lacks a concept of a session there exists no agreed upon symmetric key for the ciphering and deciphering of all packets within a session. Instead, DIDComm v2 applies the Elliptic-curve Diffie–Hellman One-Pass Unified Model (ECDH-1PU) key derivation process per message, making it more computational expensive to handle ciphered communication than using a symmetric session key for all messages of a session like in TLS. The comparably higher costs in terms of overhead and CPU usage lead to significantly higher data transfer times for DIDComm in comparison to DID Connect/TLS, especially in case a high amount of packets is transferred per session. In this TCP-based experimental setup with a reliable communication link, a TLS channel established with DID Connect therefore enables a quicker end-to-end data transfer after only few packets for payload size below 1 kB and instantly for payload sizes above 10 kB compared to a message transfer based on DIDComm. For example, 100 packets with payload size of 10 kB each is transferred 2.16 times faster via a DID Connect-established TLS channel than via DIDComm over a TCP channel.

### v2

# DID Link: Authentication in TLS with Decentralized Identifiers and Verifiable Credentials

@@ Abstract
Authentication in TLS is predominately carried out with X.509 digital certificates issued by certificate authorities (CA). The centralized nature of current public key infrastructures, however, comes along with severe risks, such as single points of failure and susceptibility to cyber-attacks, potentially undermining the security and trustworthiness of the entire system. With Decentralized Identifiers (DID) alongside distributed ledger technology, it becomes technically feasible to prove ownership of a unique identifier without requiring an attestation of the proof’s public key by a centralized and therefore vulnerable CA. This article presents DID Link, a novel authentication scheme for TLS 1.3 that empowers entities to authenticate in a TLS-compliant way with self-issued X.509 certificates that are equipped with ledger-anchored DIDs instead of CA-issued identifiers. It facilitates the exchange of tamper-proof and 3rd-party attested claims in the form of DID-bound Verifiable Credentials after the TLS handshake to complete the authentication with a full identification of the communication partner. A prototypical implementation shows comparable TLS handshake durations of DID Link if verification material is cached and reasonable prolongations if it is obtained from a ledger. The significant speed improvement of the resulting TLS channel over a widely used, DID-based alternative transport protocol on the application layer demonstrates the potential of DID Link to become a viable solution for the establishment of secure and trustful end-to-end communication links with decentrally managed digital identities.
@@ I Introduction
This article presents DID Link, a novel authentication scheme for TLS that empowers both parties to alternatively authenticate either with pseudo-anonymous DIDs or in combination with VCs instead of using CA-issued X.509 certificates or raw public keys. With DID Link, self-issued X.509 certificates are used by the involved parties as a container to communicate their DIDs and associated verification material without violating the X.509v3 specification. This makes it possible with DID Link to rely on the established way parties are authenticated during the TLS handshake by only - in addition - integrating a custom certificate validation process tailored to DID-equipped X.509 certificates. Atop the pseudo-anonymous authentication during the TLS handshake with DIDs, in DID Link, the parties can additionally exchange DID-bound VCs afterwards, to complete the authentication scheme with a clear identification of subjects that are referred to by the pseudo-anonymous DIDs used during the TLS handshake. Since DID Link remains fully compliant with TLS 1.3, the parties establishing a TLS connection with DID Link benefit from the widespread support, level of standardization, and thoroughly investigated security of TLS with X.509-based authentication and, at the same time, take advantage of the key properties of decentralized identity management, namely increased data privacy and protection, improved robustness against failures, and enhanced security due to a lack of a single point of vulnerability.
@@ V Concept
With DID Link, each entity - the client and the server - is assumed to be equipped with its own DID. The corresponding DID docs are anchored in a commonly operated and governed distributed ledger, accessible by all entities. The pseudo-anonymous authentication of an entity occurs within the TLS 1.3 handshake. Subsequently, the identification of the subject takes place at the new identification sub layer via the exchange of VCs immediately following the successful completion of the TLS handshake. The identification step is optional because both parties might know each other and each others’ DIDs from past interactions. Similar to the approach presented in [17], self-signed X.509 certificates are used in DID Link during the TLS handshake as a standardized means to communicate the identifier as well as the public key of the certificate’s subject. In contrast to [17], a W3C DID is used as an identifier. It is stored in the SAN extension instead of the subject field, as sketched in [19], due to the strict naming rules of the X.509 subject field and the 64 byte limit of the common name (CN) attribute. The public key of the self-signed X.509 certificate corresponds to a public key of the subject’s DID doc in the distributed ledger. An exemplary message exchange of DID Link, comprising the pseudo-anonymous authentication with DIDs in the TLS handshake and the identification with VCs at the application layer, is illustrated in Fig. 4.
@@
Figure 4: Using DIDs in X.509 certificates during the TLS handshake for the pseudo-anonymous authentication and VPs afterwards to (mutually) identify entities after the TCP connection establishment with DID Link.
@@
As described in Section II, the client and server verify each others certificate, the CertificateVerify and the Finished messages. With a CA-issued X.509 certificate, the integrity needs to be tested by verifying the digital signature of the CA. Moreover, if intermediary CAs are present then the chain of certificates needs to be verified all way up to the root CA. However, if a self-signed X.509 certificate’s subject refers to a DID, the certificate validation becomes significantly simpler. The integrity of a self-signed and DID-based X.509 certificate is a prerequisite, but does not get explicitly tested through the verification of the certificate’s digital signature. DID Link uses the certificate only as a vehicle to communicate the binding. However, the validity of the binding between the given DID in the SAN field and the public key in the Subject Public Key Info field needs to be checked with other means. Since the distributed ledger guarantees the uniqueness of a DID and that only the DID owner is able to alter the corresponding DID doc, it can, in the role of an independent technical trust anchor, confirm or deny the validity of a binding between a DID and a public key. If a binding’s validity is confirmed by or through the distributed ledger, the verifying entity still does not know whether the other entity is the DID owner or an entity just pretending to own the DID. The DID ownership proof is missing. Since the CertificateVerify message is signed with the private key corresponding to the public key of the certificate, it serves as a valid DID ownership proof. Technically. the self-signed X.509 certificate would actually only need to contain the DID, since the public key could be queried by resolving the DID doc from the ledger. However, providing the public key up front within the certificate speeds up the validation process because the verification of the CertificateVerify message can be carried out in parallel to the binding check via the ledger. The order of validation checks is therefore irrelevant. If the binding and the CertifyVerify message are both valid, the pseudo-anonymous authentication with a DID is completed.
@@
With DID Link, the entities are free to choose a suitable form of authentication based on the scenario-specific requirements, the presence of CA-issued X.509 certificates and/or DIDs/VCs, and potentially prior knowledge about each other. DID Link supports both, one-way and mutual authentication. It gives entities the opportunity to fall back on CA-issued X.509 certificates for the authentication in case self-signed X.509 certificates with DIDs are not supported. An entity can either authenticate with a CA-issued X.509 certificate, a self-issued and DID-based X.509 certificate only, or in combination with a VC. In hybrid mode, one party can make use of a self-issued and DID-based X.509 certificate or in combination with a VC to authenticate itself while the other party falls back on a CA-issued X.509 certificate. If the client does not demand the server to use a specific identifier for the authentication (by using the SNI extension), the server can provide a default CA-based X.509 certificate (Cerdef) or a default DID (DIDdef) within a self-issued X.509 certificate. Table I lists the numerous authentication options the entities can choose from together with the extensions present in the ClientHello and ServerHello messages. Since the extensions CPP and SPA consistently co-occur whenever at least one party expresses an intention to showcase VCs, they have been excluded from the table for the sake of readability.
@@ VI Evaluation
DID Link represents an alternative authentication scheme for TLS 1.3, bridging the gap from authentication with centrally-approved to ledger-approved identification material. The practicability of DID Link depends upon how it performs in terms of speed against the conventional way of authentication in TLS with CA-issued X.509 certificates. However, in contrast to the legacy authentication scheme, DID Link comes in two flavors: a pseudo-anonymous authentication with DIDs only or an identification with VCs in addition. A meaningful assessment of its performance is therefore only possible by comparing both variants independently with their contemporary counterparts. The practical usefulness of DID Link also hinges on the bandwidth utilization of the resulting TLS channel. Hence, a comparison against an alternative data transport solution that also employs DIDs for authentication purposes is required to properly evaluate the proposed solution in terms of efficiency.
@@ VI-A Methodology
From a conceptual point of view, a self-issued X.509 certificate with a subject of choice in the subject field is the closest analog to a self-issued x.509 certificate with a DID in the ASN field. In both cases, the certificate owner can prove ownership of the contained public key. Both types of certificates lack a proof of the binding’s validity by a 3rd party. However, with VDR-anchored DIDs, this proof is provided by the VDR. With legacy self-issued X.509 certificates, there is no such standardized and trustful out-of-bound option to check the binding’s validity. Hence, its validity can’t be tested unless the identifier results from the key itself, e.g., through hashing as applied in the Bitcoin network to derive the wallet address from the public key. A derived identifier is of pseudo-anonymous nature like a DID. It does not share the property of being resolvable but its uniqueness is guaranteed probabilistically in the same way as with DIDs. The pseudo-anonymous authentication scheme with DIDs in DID Link was therefore evaluated by testing its performance against the authentication by means of a self-issued X.509 certificate with a derived identifier. If the pseudo-anonymous authentication is supplemented by an identification with VCs afterwards then DID Link accomplishes the same goal as an authentication with legacy X.509 certificates issued by CAs, namely proving to own a 3rd-party issued identifier. First, the DID ownership is proven during the TLS handshake with self-issued and DID-equipped X.509 certificates. Afterwards, the subject referred to by the DID is revealed with a DID-bound VC that contains the 3rd party attested identifier and optionally additional identity claims of the subject. The combination of a pseudo-anonymous authentication and an identification with DID Link is therefore evaluated by comparing it as a whole against the conventional authentication with CA-issued X.509 certificates.
@@
For the evaluation of the bandwidth utilization, a TLS channel established with DID Link was compared against a communication link that uses DIDComm [27] instead. DIDComm is an application layer transport protocol built atop the concept of DIDs. It enables two parties to exchange messages in form of JSON Web Encryption (JWE) envelopes that are encrypted using the verification material found within each others’ DID docs, ensuring end-to-end confidentiality, integrity, and authenticity. DIDComm operates without the need for established connections and is agnostic to the transport layer protocol, with implementations available across various programming languages. Despite being connectionless and not being a transport layer protocol, it was selected for the comparison due to its widespread use, the ongoing standardization by the DIF, and the lack of an available DID-enabled transport layer protocol.
@@ VI-B Setup
The proof of concept, comprising a client and a server, was implemented with Python 3.10 using OpenSSL 3.0.2 wrapped with pyOpenSSL 23.3. The VC presentation protocol was built from scratch, in compliance with the DIF Presentation Exchange 2.X.X specification. The VCs were encoded as SD-JWTs because of its simplicity compared to JSON-LD and other formats. For the performance comparison, the prototype relies on Veramo’s implementation of DIDComm v2444https://veramo.io/. The distributed ledgers used for the experiments were powered by Hyperledger Indy [^5] https://www.hyperledger.org/projects/hyperledger-indy. During the experiments, DID Link was tested with different ledger realizations. A locally deployed ledger based on the VON Network project [^6] https://github.com/bcgov/von-network was used to simulate a setup in which the client and server are each assumed to be co-located with a node of the ledger, making a ledger access practically a local rather than a remote operation. To simulate remote ledger access, two publicly accessible ledgers were used: the test ledger of the BCovrin Test Indy Network operated by the Province of British Columbia [^7] http://test.bcovrin.vonx.io/ and the test ledger of the German IDunion project [^8] https://idunion.org/. The latter is commonly governed and operated by the IDunion project members, coming from various economic sectors such finance, transportation, telecommunication, and manufacturing. In the experimental setup, the client was operated on a machine with an Intel Core i5-6200U CPU with 2.30 GHz and 8 GB RAM while the server resided on a machine with an Intel Core i7-4790K CPU with 4.00 GHz and 32 GB RAM. Ubuntu-based KDE Neon 5.27 served as the operating system on both machines. They were located 7 km apart with an average ping time of 12.61ms and a bandwidth of 52 Mb/s (download) and 11 Mb/s (upload) of the client and respectively 57 Mb/s (download) and 22 Mb/s (upload) of the server. Access to the Internet was provisioned by Wi-Fi access points on both sites.
@@ VI-C Results
The results show that if VDR-anchored DIDs are cached at the client and server (config. V) or peer DIDs are used instead (config. VI), the TLS handshake with DID Link performs similar in terms of speed as their legacy counterpart (config. I). As soon as a DID doc resolution via a VDR is required, the TLS handshake with DID Link takes significantly longer. For example, a resolution via a locally deployed VON ledger at the server in config. VIII increases the total handshake time in comparison to config. VI by a factor of 3.23 while it increases by factor of 31.58 with the BCoverin ledger in config. X and 12.12 with the IDunion ledger in config. XII. The reasons for the high resolution times via a remote ledger could be manifold, ranging from the implementation of the resolution logic and the ledger itself, over the communication latency with the ledger and the amount of transactions in the ledger, to the genesis file that specifies the settings for the ledger access. The prototype was operated in Germany, with a low latency towards the Germany-based IDunion ledger and a higher latency towards the BCoverin ledger operated in western Canada. The BCoverin ledger persisted approx. 640k transactions while the IDunion test ledger contained only approx. 117k transactions. During the evaluation, the Indy CLI driver [^9] https://github.com/hyperledger/indy-cli-rs provided by the Hyperledger Indy project was used to resolve the DID docs as it turned out in prior experiments with a locally deployed VON ledger that the resolution via a locally deployed Universal Resolver [^10] https://dev.uniresolver.io/ or directly with the Indy driver of the Universal Resolver is on average 2.03 respectively 1.89 times slower than a resolution with the Indy CLI. A genesis file points to a set of ledger nodes usable for the resolution. This selection is of utmost importance since slow or non-reachable ledger nodes in the genesis file negatively impact the overall resolution performance, as experienced with different genesis file versions for the IDunion ledger.
@@
In a second experiment, the durations of the mutual identification step after the TLS handshake were measured in addition for the configs. VI to XII. Outliers (approx. 1%) were removed since non-reachable ledger nodes during the DID resolution led to timeouts that significantly skewed the average durations. Since the pseudo-anonymous authentication together with the identification step of DID Link reassembles an authentication with a CA-issued X.509 certificate, the durations for the latter were measured in configs. II to IV for comparison reasons. The messages exchanged by the presentation protocol during the identification in the proof of concept differs slightly to the sequence of messages sketched schematically in Fig. 4. The DIF Presentation Exchange specification demands that a message containing the VP is the response to a verifier’s request to present a VP. Hence, in the prototype, each party triggers the other party’s presentation of a VP by requesting it explicitly. However, the mutual VP presentations in the prototype were conducted in parallel to improve the identification step’s performance.
@@
The results show that when compared with an authentication based on CA-issued X.509 certificates, it takes at the client about 3.19 and 6.43 times longer to identify a subject with DID Link in configs. VI and VII and about 42.94 and 34.86 times longer if a remote ledger is queried for the DID docs in configs. X and XII. The times to resolve the DID docs of the client, server, and the issuers of the VCs dominate the overall time of the pseudo-anonymous authentication and identification steps of DID Link, making the DID resolution the major bottleneck of the proposed solution. In all configs. from VI to XII, it seems as it takes significantly longer for the client to identify the server than vice versa. This phenomenon results from the fact that the client has to wait for the server to process the client’s Certificate, CertificateVerify, and Finished messages before the server can proceed with the identification step. Since the client emits the first identification message, namely the VC presentation request, immediately after emitting the Finished message, it is already engaged in the identification while the server is still busy in the TLS handshake. However, an identification step with VCs is assumed to happen only once two entities have to become known to each other. Once this is done, a pseudo-anonymous authentication with DIDs is sufficient since the subjects the DIDs refer to are known. If commonly agreed upon DID doc update policies are established upon all participants of a ledger, DID docs might only change at regular or defined points in time so that most of the time locally cached versions of DID docs are sufficient for a secure TLS connection establishment. A pseudo-anonymous authentication in config. VI with DID Link is by a factor of 1.05 even slightly faster compared to the authentication in config. I. In contrast to DID Link and due to tight coupling of a CA-issued identifier and the corresponding key material in a signed artifact, the authentication with CA-issued X.509 certificates needs to be repeated at every TLS connection establishment.
@@
TABLE II: Kilobytes transferred with DID Link ($\Diamond$) or DIDComm ($\blacklozenge$) depending on number of packets per session and payload size
Finally, the performances of TLS 1.3 (with DID Link) was compared with DIDComm version 2 in terms of bandwidth utilization and the average time to transfer data in the experimental setup. For a fair comparison, DIDComm applied TCP as the transport layer protocol. The total bytes transferred by the client towards the server were measured for different payload sizes and number of packets per session. It includes two session tickets for TLS with DID Link of total size 1.5 kB for a potential session resumption. The results are shown in Table II. For the data transfer time measurements, the client established a TLS connection (with DID Link) to the server and transferred a certain amount of packets via the same session. A TCP session was established instead of TLS for the measurements with the DIDComm protocol. The average time for the client to send a bulk of packets to the server in the experimental setup depending on the amount of packets per bulk after 50 runs are shown in Fig. 6.
For payload sizes up until 1 kB, the initial protocol overhead with DID Link for the transfer of a single packet is significantly higher than the one experienced with DIDComm because of the required TLS handshake. However, once the session is established, the protocol overhead with DID Link/ TLS becomes smaller in comparison to DIDComm. The point at which a connection with DIDConnect utilizes the channel better depends on the payload size. It needs 5 packets with payload size of 1 to 100 bytes for a TLS connection to be more efficient than DIDComm or 4 packets for payload size 1 kB. Since DIDComm lacks a concept of a session there exists no agreed upon symmetric key for the ciphering and deciphering of all packets within a session. Instead, DIDComm v2 applies the Elliptic-curve Diffie–Hellman One-Pass Unified Model (ECDH-1PU) key derivation process per message, making it more computational expensive to handle ciphered communication than using a symmetric session key for all messages of a session like in TLS. The comparably higher costs in terms of overhead and CPU usage lead to significantly higher data transfer times for DIDComm in comparison to DID Link/TLS, especially in case a high amount of packets is transferred per session. In this TCP-based experimental setup with a reliable communication link, a TLS channel established with DID Link therefore enables a quicker end-to-end data transfer after only few packets for payload size below 1 kB and instantly for payload sizes above 10 kB compared to a message transfer based on DIDComm. For example, 100 packets with payload size of 10 kB each is transferred 2.16 times faster via a DID Link-established TLS channel than via DIDComm over a TCP channel.